OpenStudy (anonymous):

Please help me i really need it!!! Which is approximately equal to the area of the circle? Use 3.14 for PI.

OpenStudy (anonymous):

Assuming radius of 1: pi * r^2 = 3.14 * 1^2 = 3.14

OpenStudy (anonymous):

If the radius is a multiple of 7, then use 22/7. If the radius is a decimal number, or it is not a multiple of 7, then use 3.14. Area of a circle= pi√r.

OpenStudy (anonymous):

ok and what is that number

OpenStudy (anonymous):

I don't know, you tell me.

OpenStudy (anonymous):

can you please help me look at the picture i posted

OpenStudy (anonymous):

What options do they give you?

OpenStudy (anonymous):

50.25

OpenStudy (anonymous):

25.12 16

OpenStudy (anonymous):

12.56

OpenStudy (anonymous):

Well, is 4 a multiple of 7?

OpenStudy (anonymous):

It isn't. :P so just use pi√r with your answer choices until you get the right answer. It helps you check.

OpenStudy (anonymous):

so 3.14 / 4

OpenStudy (anonymous):

ok i got it thanks so much!!!!1`

OpenStudy (anonymous):

Welcome. :)